Archive for the ‘Issues’ Category

Downtime Nov. 11th 2011

Today Nov. 11th, 2011, at 5:34 PM Panama time a our facility experienced a power failure and the generator had problems starting. At that time the electric team started working and the problem was fixed. The total downtime was 01hr 35min.

All nodes are back to normal. If you are experiencing problems please send us an email to with the details of your service so we can verify it for you.

Jorge Moreno
NOC Manager


Downtime Feb 10th 2011

Today starting at 7:00 PM -5 GMT we start receiving massive traffic in one of our upstream providers. Usually this kind of attack is automatically detected and mitigated but today for some unknown reason the mitigation equipment was not detecting or telling us nothing about the traffic. After long troubleshooting with the manufacturer we were able to identify and mitigate the attack. At this time everything is running fine and we do not expect to have further issue with this case. We apologize for any inconvenience this may cause you.


Jorge Moreno
Manager at CCI

Network Intermittency June 1st 2010


At around 11 AM EST the Edge Router facing one of our carriers started failing possible do to a network based attack.
Our network team with help of the carrier were able to solve the intermittency at around 1 PM EST. We are monitoring the router and traffic. At this point all traffic is back to usual.

Jorge Moreno
Manager at CCI



Today we continue experimenting network problems. Fortunately we were able to find the root of the problem and it is now totally solved. We do not expect any further outage regarding this issue.

Jorge Moreno

Shared-5 Hosting node maintenance

Apirl 6th 9:00 PM EST UPDATE:

The new server is working as expected. The majority of the clients are now imported and working. If your website is not running or is presenting issues please contact us at with your name, domain name and issue so we can assist you.

April 6th 3:00 P.M. EST UPDATE:

Web traffic is now disabled. We are currently doing clients backup through Plesk backup tools. The new server is under setup and testing and soon we will start moving backups. Advise that this is a time consuming task. We are doing our best efforts to keep this as short as possible. Email traffic is still active.

April 6th 12:43 P.M. EST UPDATE:

Unfortunatelly the bit per bit restore of shared-5 server failed because there are some damaged sectors in both hard disk. For this reason we are going to schedule another maintenance for tonight to try other strategy.

Advise that this maintenance is preventive but very important to be completed. All clients data is safe and we will be reactivating web traffic in order to bring all services up and running.


Starting  april 5th at 10:00 P.M. EST shared-5 hosting node is under an urgency maintenance due to  hard disk malfunction. By that time we started to do bit per bit migration to a new and more powerful server. By the time we wrote this update, we are at 10% of the restore of all clients data to the new server and we expect to be finished at 4:00 P.M. april 6th.

As the migration is going to take time we have activated the email service so that your e-mail communication is not affected. The web traffic to the websites won’t be activated in order to keep data integrity after migration is ended.

We appreciate your understanding and be confident that our team is working around the clock in this issue. If you have any question please email us at


Jorge Moreno
NOC Manager
Cyber Cast Intl

Partial Network Downtime

Today at around 3:00 A.M. Panama time our atlantic route via columbus network failed. We are still checking what caused the issue. All the affected traffic was moved using other route. At this point everything is fine. We will continue monitoring. Sorry for the inconveniences.

Windows VPS problems

Windows node AlphaNode is experiencing issues with virtuzzo engine file corruption . This is causing erratic behavior of the main node. In order to fix the issue we will be taking backup of each virtual machine and we are going to rebuild the server. We estimate to have a fresh system by tomorrow Oct 30th..

I’ll update this ticket once we have Virtuozzo is back up.

Parent DNS server for .pa domains down

Update: The issue is now resolved.

Since morning Aug. 17th the parent DNS servers for domains .pa are not responding. This is out of our control as this servers are hosted by the “Universidad Tecnologica de Panama”. We have reported the issue to them and they are aware.

During this issue all are not going to be reached.

Any further question contact us via live chat or at

DDoS mitigation Services expanded


We are pleased to announce that we have fully integrated new DDoS mitigation techniques. This includes the upgrade of the existing  mitigation network, new  servers with high performance http servers capable of handle SYN flood and Get attacks. All this from server physcially lcoated in Panama.  No other Panama hosting firm will offer the uptime we offer under DDoS Attack.

Alpha Windows VPS Problems

UPDATE (Aug. 6, 2009): Probelm solved. Finally Parallels team was able to solve the problem.  At this time ALL VPS are running fine. If you still have problems please let us know by sending an email to

UPDATE (Aug. 2, 2009): First, we want to thank you all for your patience with this issue. At this point Parallels team and  CCI Team are working around the clock to solve the server problem which is related to a corruptions of system files that are causing Virtuozzo vitalization software to crash.  We will keep you updated.


Since 2 days ago our Alpha Host which offers Virtuozzo VPS services is having problems. At this point we are working with our partner Parallels dealing with the problem. We hope to be able to bring it up very soon.  We also want to let you know that the files and VPS are there with no problem as the problem is related to the main windows platform.

As an option we can deliver fresh installations of your VPS packages in a newer server. We are also compensating all affected clients crediting 20% in the next invoice.

If you need further information please contact us at or at live chat.

Electrical Issue on Racks 6 and 7

Today at around 6:10 AM EST Rack 6 and 7 had electrical issues due to a alerted UPS module. The problem was solved at around 6:30 AM EST. We will require to do a maintenance soon to satisfy our quality levels. We will be announcing the specific date. This only affects a few dedicated servers. If you need any further information please contact us at or at live chat.

UPDATE ( july 27th 2009 2:30 P.M. EST):  We have successfully changed the UPS module. The unit is back to its normal operation. There is no scheduled maintenance for this unit.